# Demonstration of DB-GPT: Next Generation Data Interaction System Empowered by Large Language Models

### v1

@@ Abstract.
The recent breakthroughs in large language models (LLMs) are positioned to transition many areas of software. The technologies of interacting with data particularly have an important entanglement with LLMs as efficient and intuitive data interactions are paramount. In this paper, we present DB-GPT, a revolutionary and product-ready Python library that integrates LLMs into traditional data interaction tasks to enhance user experience and accessibility. DB-GPT is designed to understand data interaction tasks described by natural language and provide context-aware responses powered by LLMs, making it an indispensable tool for users ranging from novice to expert. Its system design supports deployment across local, distributed, and cloud environments. Beyond handling basic data interaction tasks like Text-to-SQL with LLMs, it can handle complex tasks like generative data analysis through a Multi-Agents framework and the Agentic Workflow Expression Language (AWEL). The Service-oriented Multi-model Management Framework (SMMF) ensures data privacy and security, enabling users to employ DB-GPT with private LLMs. Additionally, DB-GPT offers a series of product-ready features designed to enable users to integrate DB-GPT within their product environments easily. The code of DB-GPT is available at Github111https://github.com/eosphoros-ai/DB-GPT which already has over 10.7k stars.
@@ 4. Conclusion
Future research directions include: 1) introducing powerful agents providing more powerful abilities, such as time series predictions (Jin et al., 2023; Xue et al., 2021, 2022b, 2023b; Shi et al., 2023) based on historical data and predictive decision abilities (Xue et al., 2022a; Qu et al., 2023); 2) the integration of more model training techniques. In addition to pre-training, the community is also interested in continual learning techniques for language models, such as continual pre-training (Jiang et al., 2023), prompt learning (Wang et al., 2022; Xue et al., 2023c).

### v3

@@ Abstract.
The recent breakthroughs in large language models (LLMs) are positioned to transition many areas of software. The technologies of interacting with data particularly have an important entanglement with LLMs as efficient and intuitive data interactions are paramount. In this paper, we present DB-GPT, a revolutionary and product-ready Python library that integrates LLMs into traditional data interaction tasks to enhance user experience and accessibility. DB-GPT is designed to understand data interaction tasks described by natural language and provide context-aware responses powered by LLMs, making it an indispensable tool for users ranging from novice to expert. Its system design supports deployment across local, distributed, and cloud environments. Beyond handling basic data interaction tasks like Text-to-SQL with LLMs, it can handle complex tasks like generative data analysis through a Multi-Agents framework and the Agentic Workflow Expression Language (AWEL). The Service-oriented Multi-model Management Framework (SMMF) ensures data privacy and security, enabling users to employ DB-GPT with private LLMs. Additionally, DB-GPT offers a series of product-ready features designed to enable users to integrate DB-GPT within their product environments easily. The code of DB-GPT is available at Github111https://github.com/eosphoros-ai/DB-GPT which already has over 10.7k stars. Please install DB-GPT for your own usage with the instructions 222https://github.com/eosphoros-ai/DB-GPT#installand watch a 5-minute introduction video on Youtube 333https://youtu.be/n_8RI1ENyl4 to further investigate DB-GPT.
@@ 4. Conclusion
Future research directions include: 1) introducing powerful agents providing more powerful abilities, such as time series predictions (Jin et al., 2023; Xue et al., 2021, 2022b, 2023b; Shi et al., 2023) based on historical data and predictive decision abilities (Xue et al., 2022a; Qu et al., 2023) and automatic data preparation (Qi et al., 2023, 2024; Qi and Wang, 2024); 2) the integration of more model training techniques. In addition to pre-training, the community is also interested in continual learning techniques for language models, such as continual pre-training (Jiang et al., 2023), prompt learning (Wang et al., 2022; Xue et al., 2023c).